# Rational Gluing in Edge Replacement Systems

### v1

@@ 3. Rationality of the Gluing Relation
Note that if two words $x_{0}\dots x_{m}$ and $y_{0}\dots y_{m}$ in $E_{\mathcal{R}}$ are the same except for $x_{m}\neq y_{m}$, then they represent adjacent edges in the $m$-the full expansion $E_{m}$ if and only if $x_{m}$ and $y_{m}$ are adjacent in $\Gamma_{\mathrm{Col}(x_{m-1})}=\Gamma_{\mathrm{Col}(y_{m-1})}$. Thus, even if by definition Type 1 transitions depend on edge adjacencies in the base or the replacement graphs, they really determine adjacencies in $E_{m}$. This is because, using distinguished colors for loops and defining their replacement systems to have a sole boundary vertex (differently from [BF19]) allows to see the expanded subgraph as an embedded copy of the replacement graph.
@@
    The run of $(x_{0},y_{0})(x_{1},y_{1})\dots$ in $\mathrm{Gl}_{\mathcal{R}}$ features exactly one Type 1 transition taking place when processing the first $(x_{m},y_{m})$ such that $x_{m}\neq y_{m}$. Moreover, $x_{0}\dots x_{m}$ and $y_{0}\dots y_{m}$ are adjacent edges in $E_{m}$ and the state of $\mathrm{Gl}_{\mathcal{R}}$ after processing $(x_{m},y_{m})$ is $q_{1}\big{(}\begin{smallmatrix}\mathrm{col}(x_{m})&\alpha\\ \mathrm{col}(y_{m})&\beta\end{smallmatrix}\big{)}$, which determines the type of adjacency by the entries $\alpha$ and $\beta$, as describe in Table 1.
@@
By Lemma 3.1, $x_{0}\dots x_{m-1}=y_{0}\dots y_{m-1}$ belongs to $E_{\mathcal{R}}$, so it is an edge of $E_{m-1}$. By the definition of a Type 1 transition, it follows that $x_{0}\dots x_{m}$ and $y_{0}\dots y_{m}$ are adjacent in $E_{m}$ and that the state of $\mathrm{Gl}_{\mathcal{R}}$ after processing $(x_{m},y_{m})$ is $q_{1}\big{(}\begin{smallmatrix}\mathrm{col}(x_{m})&\alpha\\ \mathrm{col}(y_{m})&\beta\end{smallmatrix}\big{)}$, where $\alpha$ and $\beta$ describe the type adjacency as expressed in Table 1. ∎
@@
Moreover, the type of adjacency between $x_{m}$ and $y_{m}$ in $\Gamma_{\mathrm{Col}(x_{m-1})}=\Gamma_{\mathrm{Col}(y_{m-1})}$ and between $x_{0}\dots x_{m}$ and $y_{0}\dots y_{m}$ in $E_{m}$ is the same, i.e., the last state of the run $q_{1}\big{(}\begin{smallmatrix}\mathrm{col}(x_{m})&\alpha\\ \mathrm{col}(y_{m})&\beta\end{smallmatrix}\big{)}$ determines the edge adjacency of $x_{0}\dots x_{m}$ and $y_{0}\dots y_{m}$ in $E_{m}$ as described in Table 1.
@@
Given two words $x_{0}\dots x_{m}$ and $y_{0}\dots y_{m}$ in $E_{\mathcal{R}}$ that only differ in $x_{m}\neq y_{m}$ and that represent adjacent edges of $E_{m}$, by Lemma 3.1 the sub-automaton $\mathrm{T}_{\mathcal{R}}$ recognizes $(x_{0},y_{0})(x_{1},y_{1})\ldots(x_{m-1},y_{m-1})$. By definition, a Type 1 transition occurs when processing the letter $(x_{m},y_{m})$, ending in a state $q_{1}\big{(}\begin{smallmatrix}\mathrm{col}(x_{m})&\alpha\\ \mathrm{col}(y_{m})&\beta\end{smallmatrix}\big{)}$ that determines the edge adjacency between $x_{m}$ and $y_{m}$ in the replacement graph $\Gamma_{\mathrm{col}(x_{m-1})}=\Gamma_{\mathrm{col}(y_{m-1})}$, and thus in $E_{m}$, as expressed in Table 1. ∎
@@
First, suppose that the transition involving $(x_{k},y_{k})$ is the first Type 2 transition, so the previous transition is of Type 1. Then, by Lemma 3.2, the edges $x_{1}\dots x_{k-1}$ and $y_{1}\dots y_{k-1}$ are adjacent and the state $q_{1}\big{(}\begin{smallmatrix}\mathrm{col}(x_{k-1})&\alpha\\ \mathrm{col}(y_{k-1})&\beta\end{smallmatrix}\big{)}$ after processing $(x_{k-1},y_{k-1})$ describes the type of adjacency as expressed in Table 1. Thoroughly going through Table 2 reveals that the column and the row of any non empty entry correspond to a pair of adjacent edges in the graph obtained by gluing the replacement graphs $\Gamma_{\mathrm{col}(x_{k-1})}$ and $\Gamma_{\mathrm{col}(y_{k-1})}$ as dictated by $\alpha$ and $\beta$, which appears as a subgraph of $E_{k}$ that contains the edges $x_{1}\dots x_{k}$ and $y_{1}\dots y_{k}$. Moreover, the resulting state $q_{1}\big{(}\begin{smallmatrix}\mathrm{col}(x_{k})&\alpha^{\prime}\\ \mathrm{col}(y_{k})&\beta^{\prime}\end{smallmatrix}\big{)}$ describes the adjacency between $x_{0}\dots x_{k}$ and $y_{0}\dots y_{k}$.
Now suppose that the transition involving $(x_{k},y_{k})$ is not the first Type 2 transition. The argument of the previous paragraph applies as soon as the last state in the run of $(x_{1},y_{1})\dots(x_{k-1},y_{k-1})$ described the adjacency of $(x_{k-1},y_{k-1})$. Thus, we can iteratively apply the argument to see that $x_{0}\dots x_{k}$ and $y_{0}\dots y_{k}$ are adjacent and that the last state of the run determines their type of adjacency. ∎
@@
First, suppose that the state after having processed $(x_{0},y_{0}),\dots(x_{m},y_{m})$ is $q_{1}\big{(}\begin{smallmatrix}\mathrm{col}(x_{m})&\alpha\\ \mathrm{col}(y_{m})&\beta\end{smallmatrix}\big{)}$ and that it describes the type adjacency as expressed in Table 1. Then a careful look at Table 2 shows that the gluing automaton can process the next letter $(x_{m+1},y_{m+1})$ with a Type 2 transition that ends in the state $q_{1}\big{(}\begin{smallmatrix}\mathrm{col}(x_{m+1})&\alpha^{\prime}\\ \mathrm{col}(y_{m+1})&\beta^{\prime}\end{smallmatrix}\big{)}$ which describes the adjacency between $x_{0}\dots x_{m+1}$ and $y_{0}\dots y_{m+1}$.
Now, by Lemma 3.3 the common prefix $x_{0}\dots x_{p}=y_{0}\dots y_{p}$ of $x_{0}\dots x_{m}$ and $y_{0}\dots y_{m}$ (where $10\leq p\leq m-$ 1) is processed by the sub-automaton $\mathrm{T}_{\mathcal{R}}$, after which a Type 1 transition brings to the state $q_{1}\big{(}\begin{smallmatrix}\mathrm{col}(x_{p+1})&\alpha\\ \mathrm{col}(y_{p+1})&\beta\end{smallmatrix}\big{)}$ that describes the adjacency of $x_{0}\dots x_{m}$ and $y_{0}\dots y_{m}$ in $E_{m}$. Reasoning as in the previous paragraph shows that this is followed by a Type 2 transition to the state $q_{1}\big{(}\begin{smallmatrix}\mathrm{col}(x_{p+2})&\alpha^{\prime}\\ \mathrm{col}(y_{p+2})&\beta^{\prime}\end{smallmatrix}\big{)}$ that describes the adjacency between $x_{0}\dots x_{p+2}$ and $y_{0}\dots y_{p+2}$. Since each new transition brings to a state that describes edge adjacency, the previous paragraph can be iteratively applied to every subsequent prefix of $x_{0}\dots x_{m+1}$ and $y_{0}\dots y_{m+1}$, concluding with a Type 2 transition from $(x_{0},y_{0})\dots(x_{m},y_{m})$ to $(x_{0},y_{0})\dots(x_{m+1},y_{m+1})$ that ends in a state that described the edge adjacency between $x_{0}\dots x_{m+1}$ and $y_{0}\dots y_{m+1}$. ∎
@@
Suppose $v=x_{0}x_{1}\dots$ and $w=y_{0}y_{1}\dots$ are equivalent. If $v=w$ then $\mathrm{Gl}_{\mathcal{R}}$ recognizes the sequence $(v,w)$ by Lemma 3.1, so we are done. Suppose that instead $v\neq w$ and let $m\geq 0$ be the smallest natural number such that $x_{m}\neq y_{m}$. Since $v\sim w$, by Definition 1.4 there exists a unique vertex on which $x_{0}\ldots x_{k}$ and $y_{0}\ldots y_{k}$ are incident as edges of $E_{k}$, for all $k\geq m$. In particular, $x_{0}\ldots x_{m}$ and $y_{0}\ldots y_{m}$ are adjacent, so by Lemma 3.3 the prefix $(x_{0},y_{0})\ldots(x_{m},y_{m})$ is recognized by $\mathrm{Gl}_{\mathcal{R}}$ and its run ends with a Type 1 transition to the state $q_{1}\big{(}\begin{smallmatrix}\mathrm{col}(x_{m})&\alpha\\ \mathrm{col}(y_{m})&\beta\end{smallmatrix}\big{)}$, which describes the type of adjacency between $x_{1}\dots x_{m}$ and $y_{0}\dots y_{m}$ in $E_{m}$.
@@
By definition of Type 1 and 2 transitions, in the schemes above $a$ is an edge of the graph $\Gamma_{i}$. Thus, $\phi(\mathrm{Gl}_{\mathcal{R}})$ is a subgraph of the color graph, so $\phi:\mathrm{Gl}_{\mathcal{R}}\to A_{\mathcal{R}}$ is a well-defined graph morphism. Thus, if $(v,w)$ is recognized $\mathrm{Gl}_{\mathcal{R}}$ then $v$ is in the edge shift $\Omega_{\mathcal{R}}$ of the color graph. The same holds for the entry $w$, since $\mathrm{Gl}_{\mathcal{R}}$ is symmetric under switching inputs $(a,b)$ with $(b,a)$ and states $q_{i}\big{(}\begin{smallmatrix}i&\gamma\\ j&\delta\end{smallmatrix}\big{)}$ with $q_{i}\big{(}\begin{smallmatrix}j&\delta\\ i&\gamma\end{smallmatrix}\big{)}$.
We now show that $v\sim w$. If $v=w$ there is nothing to prove, so let $v\neq w$. By Lemma 3.2, the run of $(v,w)$ features finitely many Type 0 transitions followed by a Type 1 transition at $(x_{0},y_{0}),\dots(x_{m},y_{m})$, for some $m\geq 0$, that brings to a state of $Q_{1}$, and $x_{0}\dots x_{m}$ is adjacent to $y_{0}\dots y_{m}$. This shows that the prefixes of $v$ and $w$ of equal length $k\leq m+1$ represent adjacent edges. Since the transitions of the gluing automaton that start in $Q_{1}$ are of Type 2 and end in $Q_{1}$, every subsequent transition of the run of $(v,w)$ is a Type 2 transition. By Lemma 3.4, the prefixes of $v$ and $w$ of equal length $k>m+1$ are adjacent, concluding our proof. ∎
@@ Acknowledgements
The authors are thankful to Francesco Matucci for helpful advices and to Jim Belk and Bradley Forrest for kindly providing the image used in Figure 4(b).

### v2

@@ 3. Rationality of the Gluing Relation
Note that if two words $x_{0}\dots x_{m}$ and $y_{0}\dots y_{m}$ in $E_{\mathcal{R}}$ are the same except for $x_{m}\neq y_{m}$, then they represent adjacent edges in the $m$-th full expansion $E_{m}$ (Definition 1.4) if and only if $x_{m}$ and $y_{m}$ are adjacent in $\Gamma_{\mathrm{Col}(x_{m-1})}=\Gamma_{\mathrm{Col}(y_{m-1})}$. Thus, even if by definition Type 1 transitions depend on edge adjacencies in the base or the replacement graphs, they really determine adjacencies in $E_{m}$. This is because identifying the initial and terminal vertices of replacement graphs for colors associated to loops (as explained in Subsection 1.1.1) allows us to see the expanded subgraph as an embedded copy of the replacement graph.
@@
    The run of $(x_{0},y_{0})(x_{1},y_{1})\dots$ in $\mathrm{Gl}_{\mathcal{R}}$ features exactly one Type 1 transition taking place when processing the first $(x_{m},y_{m})$ such that $x_{m}\neq y_{m}$. Moreover, $x_{0}\dots x_{m}$ and $y_{0}\dots y_{m}$ are adjacent edges in $E_{m}$ and the state of $\mathrm{Gl}_{\mathcal{R}}$ after processing $(x_{m},y_{m})$ is $q_{1}\big{(}\begin{smallmatrix}\mathrm{col}(x_{m})&\alpha\\ \mathrm{col}(y_{m})&\beta\end{smallmatrix}\big{)}$, which determines the type of adjacency by the entries $\alpha$ and $\beta$, as describe in Table I.
@@
By Lemma 3.1, $x_{0}\dots x_{m-1}=y_{0}\dots y_{m-1}$ belongs to $E_{\mathcal{R}}$, so it is an edge of $E_{m-1}$. By the definition of a Type 1 transition, it follows that $x_{0}\dots x_{m}$ and $y_{0}\dots y_{m}$ are adjacent in $E_{m}$ and that the state of $\mathrm{Gl}_{\mathcal{R}}$ after processing $(x_{m},y_{m})$ is $q_{1}\big{(}\begin{smallmatrix}\mathrm{col}(x_{m})&\alpha\\ \mathrm{col}(y_{m})&\beta\end{smallmatrix}\big{)}$, where $\alpha$ and $\beta$ describe the type adjacency as expressed in Table I. ∎
@@
Moreover, the type of adjacency between $x_{m}$ and $y_{m}$ in $\Gamma_{\mathrm{Col}(x_{m-1})}=\Gamma_{\mathrm{Col}(y_{m-1})}$ and between $x_{0}\dots x_{m}$ and $y_{0}\dots y_{m}$ in $E_{m}$ is the same, i.e., the last state of the run $q_{1}\big{(}\begin{smallmatrix}\mathrm{col}(x_{m})&\alpha\\ \mathrm{col}(y_{m})&\beta\end{smallmatrix}\big{)}$ determines the edge adjacency of $x_{0}\dots x_{m}$ and $y_{0}\dots y_{m}$ in $E_{m}$ as described in Table I.
@@
Given two words $x_{0}\dots x_{m}$ and $y_{0}\dots y_{m}$ in $E_{\mathcal{R}}$ that only differ in $x_{m}\neq y_{m}$ and that represent adjacent edges of $E_{m}$, by Lemma 3.1 the sub-automaton $\mathrm{T}_{\mathcal{R}}$ recognizes $(x_{0},y_{0})(x_{1},y_{1})\ldots(x_{m-1},y_{m-1})$. By definition, a Type 1 transition occurs when processing the letter $(x_{m},y_{m})$, ending in a state $q_{1}\big{(}\begin{smallmatrix}\mathrm{col}(x_{m})&\alpha\\ \mathrm{col}(y_{m})&\beta\end{smallmatrix}\big{)}$ that determines the edge adjacency between $x_{m}$ and $y_{m}$ in the replacement graph $\Gamma_{\mathrm{col}(x_{m-1})}=\Gamma_{\mathrm{col}(y_{m-1})}$, and thus in $E_{m}$, as expressed in Table I. ∎
@@
First, suppose that the transition involving $(x_{k},y_{k})$ is the first Type 2 transition, so the previous transition is of Type 1. Then, by Lemma 3.2, the edges $x_{1}\dots x_{k-1}$ and $y_{1}\dots y_{k-1}$ are adjacent and the state $q_{1}\big{(}\begin{smallmatrix}\mathrm{col}(x_{k-1})&\alpha\\ \mathrm{col}(y_{k-1})&\beta\end{smallmatrix}\big{)}$ after processing $(x_{k-1},y_{k-1})$ describes the type of adjacency as expressed in Table I. Thoroughly going through Table II reveals that the column and the row of any non empty entry correspond to a pair of adjacent edges in the graph obtained by gluing the replacement graphs $\Gamma_{\mathrm{col}(x_{k-1})}$ and $\Gamma_{\mathrm{col}(y_{k-1})}$ as dictated by $\alpha$ and $\beta$, which appears as a subgraph of $E_{k}$ that contains the edges $x_{1}\dots x_{k}$ and $y_{1}\dots y_{k}$. Moreover, the resulting state $q_{1}\big{(}\begin{smallmatrix}\mathrm{col}(x_{k})&\alpha^{\prime}\\ \mathrm{col}(y_{k})&\beta^{\prime}\end{smallmatrix}\big{)}$ describes the adjacency between $x_{0}\dots x_{k}$ and $y_{0}\dots y_{k}$.
Now suppose that the transition involving $(x_{k},y_{k})$ is not the first Type 2 transition. The argument of the previous paragraph applies as soon as the last state in the run of $(x_{1},y_{1})\dots(x_{k-1},y_{k-1})$ describes the adjacency of $(x_{k-1},y_{k-1})$. Thus, we can iteratively apply the argument to see that $x_{0}\dots x_{k}$ and $y_{0}\dots y_{k}$ are adjacent and that the last state of the run determines their type of adjacency. ∎
@@
First, suppose that the state after having processed $(x_{0},y_{0}),\dots(x_{m},y_{m})$ is $q_{1}\big{(}\begin{smallmatrix}\mathrm{col}(x_{m})&\alpha\\ \mathrm{col}(y_{m})&\beta\end{smallmatrix}\big{)}$ and that it describes the type adjacency as expressed in Table I. Then a careful look at Table II shows that the gluing automaton can process the next letter $(x_{m+1},y_{m+1})$ with a Type 2 transition that ends in the state $q_{1}\big{(}\begin{smallmatrix}\mathrm{col}(x_{m+1})&\alpha^{\prime}\\ \mathrm{col}(y_{m+1})&\beta^{\prime}\end{smallmatrix}\big{)}$ which describes the adjacency between $x_{0}\dots x_{m+1}$ and $y_{0}\dots y_{m+1}$.
Now, by Lemma 3.3 the common prefix $x_{0}\dots x_{p}=y_{0}\dots y_{p}$ of $x_{0}\dots x_{m}$ and $y_{0}\dots y_{m}$ (where $10\leq p\leq m-$ 1) is processed by the sub-automaton $\mathrm{T}_{\mathcal{R}}$, after which a Type 1 transition brings us to the state $q_{1}\big{(}\begin{smallmatrix}\mathrm{col}(x_{p+1})&\alpha\\ \mathrm{col}(y_{p+1})&\beta\end{smallmatrix}\big{)}$ that describes the adjacency of $x_{0}\dots x_{p+1}$ and $y_{0}\dots y_{p+1}$ in $E_{p+1}$. Reasoning as in the previous paragraph shows that this is followed by a Type 2 transition to the state $q_{1}\big{(}\begin{smallmatrix}\mathrm{col}(x_{p+2})&\alpha^{\prime}\\ \mathrm{col}(y_{p+2})&\beta^{\prime}\end{smallmatrix}\big{)}$ that describes the adjacency between $x_{0}\dots x_{p+2}$ and $y_{0}\dots y_{p+2}$. Since each new transition results in a state that describes edge adjacency, the previous paragraph can be iteratively applied to every subsequent prefix of $x_{0}\dots x_{m+1}$ and $y_{0}\dots y_{m+1}$, concluding with a Type 2 transition from $(x_{0},y_{0})\dots(x_{m},y_{m})$ to $(x_{0},y_{0})\dots(x_{m+1},y_{m+1})$ that ends in a state that describes the edge adjacency between $x_{0}\dots x_{m+1}$ and $y_{0}\dots y_{m+1}$. ∎
@@
Suppose $v=x_{0}x_{1}\dots$ and $w=y_{0}y_{1}\dots$ are equivalent. If $v=w$ then $\mathrm{Gl}_{\mathcal{R}}$ recognizes the sequence $(v,w)$ by Lemma 3.1, so we are done. Suppose that instead $v\neq w$ and let $m\geq 0$ be the smallest natural number such that $x_{m}\neq y_{m}$. Since $v\sim w$, by Definition 1.7 there exists a unique vertex on which $x_{0}\ldots x_{k}$ and $y_{0}\ldots y_{k}$ are incident as edges of $E_{k}$, for all $k>m$ (for $k=m$ the edges may be parallel, in which case there may be two such vertices). In particular, $x_{0}\ldots x_{m}$ and $y_{0}\ldots y_{m}$ are adjacent, so by Lemma 3.3 the prefix $(x_{0},y_{0})\ldots(x_{m},y_{m})$ is recognized by $\mathrm{Gl}_{\mathcal{R}}$ and its run ends with a Type 1 transition to the state $q_{1}\big{(}\begin{smallmatrix}\mathrm{col}(x_{m})&\alpha\\ \mathrm{col}(y_{m})&\beta\end{smallmatrix}\big{)}$, which describes the type of adjacency between $x_{1}\dots x_{m}$ and $y_{0}\dots y_{m}$ in $E_{m}$.
@@
By definition of Type 1 and 2 transitions, in the schemes above $a$ is an edge of the graph $\Gamma_{i}$. Thus, $\phi(\mathrm{Gl}_{\mathcal{R}})$ is a subgraph of the color graph, so $\phi:\mathrm{Gl}_{\mathcal{R}}\to A_{\mathcal{R}}$ is a well-defined graph morphism. In fact, it is easy to see that $\phi(\mathrm{Gl}_{\mathcal{R}})$ is precisely $A_{\mathcal{R}}$. Thus, if $(v,w)$ is recognized by $\mathrm{Gl}_{\mathcal{R}}$ then $v$ is in the edge shift $\Omega_{\mathcal{R}}$ of the color graph. The same holds for the entry $w$, since $\mathrm{Gl}_{\mathcal{R}}$ is symmetric under switching inputs $(a,b)$ with $(b,a)$ and states $q_{i}\big{(}\begin{smallmatrix}i&\gamma\\ j&\delta\end{smallmatrix}\big{)}$ with $q_{i}\big{(}\begin{smallmatrix}j&\delta\\ i&\gamma\end{smallmatrix}\big{)}$.
We now show that $v\sim w$. If $v=w$ there is nothing to prove, so let $v\neq w$. By Lemma 3.2, the run of $(v,w)$ features finitely many Type 0 transitions followed by a Type 1 transition at $(x_{0},y_{0}),\dots(x_{m},y_{m})$, for some $m\geq 0$, that brings us to a state of $Q_{1}$, and $x_{0}\dots x_{m}$ is adjacent to $y_{0}\dots y_{m}$. This shows that the prefixes of $v$ and $w$ of equal length $k\leq m+1$ represent adjacent edges. Since the transitions of the gluing automaton that start in $Q_{1}$ are of Type 2 and end in $Q_{1}$, every subsequent transition of the run of $(v,w)$ is a Type 2 transition. By Lemma 3.4, the prefixes of $v$ and $w$ of equal length $k>m+1$ are adjacent, concluding our proof. ∎
@@ Acknowledgements
The authors are thankful to Francesco Matucci for helpful advice, to Jim Belk and Bradley Forrest for kindly providing the image used in Figure 6(b) and to the anonymous referee for insightful observations that led to clearer exposition.
[…]
[heading=bibintoc]